### v2

# Hausdorffness of certain nilpotent cohomology spaces

Abstract.
Let $(\pi,V)$ be a smooth representation of a compact Lie group $G$ on a quasi-complete locally convex complex topological vector space. We show that the Lie algebra cohomology space ${\mathrm{H}}^{\bullet}(\mathfrak{u},V)$ and the Lie algebra homology space ${\mathrm{H}}_{\bullet}(\mathfrak{u},V)$ are both Hausdorff, where $\mathfrak{u}$ is the nilpotent radical of a parabolic subalgebra of the complexified Lie algebra $\mathfrak{g}$ of $G$.
Key words and phrases:
smooth representation, nilpotent cohomology, cubic Dirac operator
2020 Mathematics Subject Classification:
22E46, 17B56
1. Introduction
Let $G$ be a Lie group. In our context, a representation of $G$ is a quasi-complete, locally convex, Hausdorff complex topological vector space $V$, together with a continuous linear action
(1.1) $G\times V\rightarrow V.$
We say that a representation $V$ of $G$ is smooth if the action map (1.1) is smooth as a map between (possibly infinite-dimensional) smooth manifolds ([3]). To emphasize the action, we will denote a representation by the pair $(\pi,V)$, where $\pi$ refers to the action.
Let $\mathfrak{g}$ be a finite-dimensional complex Lie algebra. A continuous $\mathfrak{g}$-module is a quasi-complete, locally convex, Hausdorff complex topological vector space $V$, together with a continuous Lie algebra action
(1.2) $\mathfrak{g}\times V\rightarrow V.$
For a continuous $\mathfrak{g}$-module $V$, recall that the $\mathfrak{g}$-cohomology is computed by the total complex
${\mathrm{Hom}}(\wedge^{\bullet}\mathfrak{g},V)=\wedge^{\bullet}(\mathfrak{g}^{*})\otimes V,$
where $\,{}^{*}$ indicates the dual space and $\wedge^{\bullet}$ indicates the exterior algebra. The complex ${\mathrm{Hom}}(\wedge^{\bullet}\mathfrak{g},V)$ carries a natural topology and the coboundary map $\operatorname{d}\,$ is continuous for this topology. The cohomology space ${\mathrm{H}}^{\bullet}(\mathfrak{g},V)$ inherits the natural subquotient topology.
Dually, the total complex for $\mathfrak{g}$-homology is
$\wedge^{\bullet}\mathfrak{g}\otimes V$
with its natural topology and the continuous boundary map $\partial$. The homology space ${\mathrm{H}}_{\bullet}(\mathfrak{g},V)$ again inherits the natural subquotient topology. Both ${\mathrm{H}}^{\bullet}(\mathfrak{g},V)$ and ${\mathrm{H}}_{\bullet}(\mathfrak{g},V)$ are not necessarily Hausdorff, since the images of $\operatorname{d}\,$ and $\partial$ are not necessarily closed.
[…]
Given a $(\mathfrak{g},K)$-module $M$ of $G$, a globalization of $M$ is defined to be a representation $V$ of $G$ together with a $(\mathfrak{g},K)$-module isomorphism between $M$ and the underlying $(\mathfrak{g},K)$-module of $V$.
When $M$ has finite length (in this case $M$ is called a Harish-Chandra module), it has four canonical globalizations: the minimal globalization $M^{\min}$, the Casselman-Wallach globalization $M^{\infty}$, the distribution globalization $M^{-\infty}$, and the maximal globalization $M^{\max}$. These globalizations are smooth representations of $G$ on Fréchet or dual Fréchet spaces, and they fit into a sequence of inclusions
$M\subset M^{\min}\subset M^{\infty}\subset M^{-\infty}\subset M^{\max}.$
For $\alpha\in\{\min,\infty,-\infty,\max\}$, Vogan conjectured that the $\mathfrak{u}$-cohomology ${\mathrm{H}}^{\bullet}(\mathfrak{u},M^{\alpha})$ is Hausdorff, which implies that ${\mathrm{H}}^{\bullet}(\mathfrak{u},M^{\alpha})$ is a smooth representation of $L$. For more details about Vogan’s conjecture, we refer to [11, Conjecture 10.3].
Obviously, Vogan’s conjecture is trivially true if $G$ is compact, since all Harish-Chandra modules are finite-dimensional in this case. Bratten and Corti proved in [1, 2] that Vogan’s conjecture holds when $M^{\alpha}$ is the minimal globalization or the maximal globalization. For related research on Hausdorffness of Lie algebra cohomologies, see [1, 4, 10, 13, 14].
Given an arbitrary smooth representation $V$ of a real reductive group $G$, which is naturally a continuous $\mathfrak{g}$-module, we may ask more generally whether the cohomology space ${\mathrm{H}}^{\bullet}(\mathfrak{u},V)$ is Hausdorff. When $G$ is compact, we establish a stronger result that ensures the Hausdorffness of ${\mathrm{H}}^{\bullet}(\mathfrak{u},V)$, as stated in the following theorem.
Theorem 1.1.
Suppose that $G$ is compact. Let $V$ be a smooth representation of $G$.
(a) Denote by
$\operatorname{d}:\wedge^{\bullet}(\mathfrak{u}^{*})\otimes V\rightarrow\wedge^{\bullet}(\mathfrak{u}^{*})\otimes V$
and
$\partial:\wedge^{\bullet}\mathfrak{u}\otimes V\rightarrow\wedge^{\bullet}\mathfrak{u}\otimes V$
the coboundary and boundary maps respectively, both of which are $L$-equivariant with respect to the natural actions of $L$. Then both of the inclusion maps
(1.3) ${\mathrm{Im}}\operatorname{d}\hookrightarrow{\mathrm{Ker}}\operatorname{d}$
and
(1.4) ${\mathrm{Im}}\,\partial\hookrightarrow{\mathrm{Ker}}\,\partial$
admit a degree-preserving $L$-equivariant continuous linear splitting.
(b) The Lie algebra cohomology space ${\mathrm{H}}^{\bullet}(\mathfrak{u},V)$ and the Lie algebra homology space ${\mathrm{H}}_{\bullet}(\mathfrak{u},V)$ are Hausdorff and quasi-complete. Moreover, with the natural actions of $L$, ${\mathrm{H}}_{\bullet}(\mathfrak{u},V)$ and ${\mathrm{H}}^{\bullet}(\overline{\mathfrak{u}},V)$ are smooth representations of $L$ that are isomorphic to each other.
In the above theorem, the images ${\mathrm{Im}}\operatorname{d}$ and ${\mathrm{Im}}\,\partial$, as well as the kernels ${\mathrm{Ker}}\operatorname{d}$ and ${\mathrm{Ker}}\,\partial$, are all equipped with the subspace topologies.
In Section 4, we will utilize Dirac cubic operators to construct explicit splittings for the inclusion maps (1.3) and (1.4) in Theorem 1.1.
Our results are related to Vogan’s conjecture through Hochschild-Serre spectral sequence, which is one motivation of our work. Recall that $V$ is a smooth representation of a real reductive group $G$. Note that $\mathfrak{u}\cap\mathfrak{k}$ is a nilpotent subalgebra of $\mathfrak{k}$, where $\mathfrak{k}=\mathfrak{g}^{\theta}$ is the complexified Lie algebra of $K$.
Ignoring the topology, there exists a convergent spectral sequence $\{E_{r}^{p,q}\}_{r\geq 0}\Rightarrow{\mathrm{H}}^{p+q}(\mathfrak{u},V)$ in the category of $\mathfrak{l}\cap\mathfrak{k}$-modules, which is called the Hochschild-Serre spectral sequence. The $E_{1}$-terms are given by $E_{1}^{p,q}={\mathrm{H}}^{n(p,q)}(\mathfrak{u}\cap\mathfrak{k},V)\otimes X_{p}$. Here $X_{p}$ is a finite-dimensional vector space and $n(p,q)$ is a certain integer depending on $p$ and $q$. For more details on the Hochschild-Serre spectral sequence, we refer to [7, Chapter V, Section 10].
Equipped with natural subquotient topology on $E_{r}$, the Hochschild-Serre spectral sequence suggests that one may prove Vogan’s conjecture by establishing the Hausdorffness of $E_{r}$-terms for every $r\geq 0$. Along this strategy, we expect that the total cohomology space ${\mathrm{H}}^{\bullet}(\mathfrak{u},V)$ is Hausdorff for certain smooth representations $V$ whose underlying $(\mathfrak{g},K)$-module are not necessarily of finite length. The Hausdorffness of $E_{0}$-terms is clear from the construction, and our results show that the $E_{1}$-terms are indeed Hausdorff. However, for $r\geq 2$, the Hausdorffness of $E_{r}$-terms is still unresolved. We hope to investigate this in future research.
Remark 1.2.
For some non-compact real reductive groups $G$ and some smooth representations $V$ of $G$, the cohomology space ${\mathrm{H}}^{\bullet}(\overline{\mathfrak{u}},V)$ may not be isomorphic to the homology space ${\mathrm{H}}_{\bullet}(\mathfrak{u},V)$. For instance, a counterexample is provided in [6, Section 8]: Let $G={\mathrm{SL}}_{2}(\mathbb{R})$ and $\mathfrak{q}=\mathfrak{so}_{2}\oplus\mathfrak{u}$, where $\mathfrak{u}=\mathbb{C}\begin{bmatrix}1&\sqrt{-1}\\ \sqrt{-1}&-1\end{bmatrix}$. Let $M$ be a Harish-Chandra module for $G$ that is a nontrivial extension of a highest weight module of highest weight $-2$ by a trivial module. Then the dimension of ${\mathrm{H}}^{\bullet}(\overline{\mathfrak{u}},M)$ is $3$ while the dimension of ${\mathrm{H}}_{\bullet}(\mathfrak{u},M)$ is $1$. Thus the comparison theorem (cf.  [1, Theorem 1]) implies that ${\mathrm{H}}^{\bullet}(\overline{\mathfrak{u}},M^{\min})\ncong{\mathrm{H}}_{\bullet}(\mathfrak{u},M^{\min})$.
Acknowledgment: The authors would like to thank Wei Xiao for suggesting the references [5, 6, 8] on the cubic Dirac operator. F. J. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the grant SFB-TRR 358/1 2023-491392403. B. Sun was supported in part by National Key R & D Program of China No. 2022YFA1005300 and New Cornerstone Science Foundation.
2. Smooth representations of compact Lie groups
From now on, $G$ is a (possibly disconnected) compact Lie group with complexified Lie algebra $\mathfrak{g}$, and $(\pi,V)$ is a smooth representation of $G$. Fix a $G$-invariant, non-degenerate, symmetric bilinear form $B$ on $\mathfrak{g}$ whose restriction to ${\mathrm{Lie}}(G)$ is real valued and negative definite.
Write $\widehat{G}$ for the set of isomorphism classes of irreducible unitary representations of $G$. For every $\lambda\in\widehat{G}$, fix an irreducible representation $(\pi_{\lambda},V_{\lambda})$ of class $\lambda$. Then the Casimir operator $\Omega_{\mathfrak{g}}$ with respect to $B$ acts on $V_{\lambda}$ via the scalar multiplication by a non-negative real scalar $c(\lambda)$.
[…]
The following theorem is proved by Harish-Chandra (see [12, Theorem 4.4.2.1]) under the assumption that $V$ is complete. The same proof works for general quasi-complete spaces as well.
Theorem 2.1.
For every $v\in V$, the Fourier series
(2.1) $\sum_{\lambda\in\widehat{G}}\pi(\overline{\chi}_{\lambda})v$
converges absolutely to $v$.
As a consequence of the above theorem, we have an identification
(2.2) $V=\left\{\left(v(\lambda)\right)_{\lambda}\in\prod_{\lambda\in\widehat{G}}V(\lambda)\,:\,\sum_{\lambda\in\widehat{G}}v(\lambda)\text{ is absolutely convergent }\right\}.$
Lemma 2.2.
Let $W\subseteq V$ be a subspace. Assume that for every $\lambda\in\widehat{G}$, the subspace $W(\lambda):=\pi(\overline{\chi}_{\lambda})W$ is closed in $V$ and contained in $W$. Then under the identification (2.2), the closure of $W$ in $V$ equals
$\left\{\left(w(\lambda)\right)_{\lambda}\in\prod_{\lambda\in\widehat{G}}W(\lambda)\,:\,\sum_{\lambda\in\widehat{G}}w(\lambda)\text{ is absolutely convergent }\right\}.$
Consequently,
$\overline{W}=\overline{\bigoplus_{\lambda\in\widehat{G}}W(\lambda)}.$
[…]
Proof.
We have that
$\pi(\overline{\chi}_{\lambda})(\overline{W})\subset\overline{\pi(\overline{\chi}_{\lambda})(W)}=\overline{W(\lambda)}=W(\lambda).$
This implies the lemma. ∎
As usual we write :=${\mathbb{N}}:=\{0,1,2,\dots\}$. We conclude this section by quoting two lemmas that will be used in Section 4 for proving the convergence of some Fourier series.
Lemma 2.3.
[12, Lemma 4.4.2.2] For every continuous seminorm $\lvert\,\cdot\,\rvert_{p}$ on $V$, there exists a continuous seminorm $\lvert\,\cdot\,\rvert_{q}$ on $V$ such that
(2.3) $\lvert\pi\left(\overline{\chi}_{\lambda}\right)v\rvert_{p}\leq\left(1+c\left(\lambda\right)\right)^{-m}\cdot(\dim V_{\lambda})^{2}\cdot\lvert\left(1+\Omega_{\mathfrak{g}}\right)^{m}v\rvert_{q}$
for all $m\in{\mathbb{N}}$, $v\in V$, and $\lambda\in\widehat{G}$.
Lemma 2.4.
For every $n\in{\mathbb{N}}$, there exists an $m\in{\mathbb{N}}$ such that the series
$\sum_{\lambda\in\widehat{G}}(\dim V_{\lambda})^{n}(1+c(\lambda))^{-m}$
converges absolutely.
Proof.
This lemma is a variation of [12, Lemma 4.4.2.3], whose proof also works in our context. ∎
3. Cubic Dirac operator and $\mathfrak{u}$-cohomology
In this section, we review the definition and basic properties of Kostant’s cubic Dirac operator. For more details about Dirac operators, we refer to [5, 6, 8].
We retain the notations from previous sections. In particular, $\mathfrak{g}$ is the complexified Lie algebra of a compact Lie group $G$, $\mathfrak{q}=\mathfrak{l}\oplus\mathfrak{u}$ is a parabolic subalgebra and $L=N_{G}(\mathfrak{q})$ is the Levi subgroup.
From this point onward, we assume that the bilinear form $B$ extends the Killing form on the semisimple part $\mathfrak{g}^{ss}$ of $\mathfrak{g}$.
From now on, we fix a Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{g}$ contained in $\mathfrak{l}$. Then $B$ induces on $\mathfrak{h}^{*}$ a non-degenerate bilinear form which we denote by (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). Write $\Delta_{\mathfrak{g}}=\Delta(\mathfrak{g},\mathfrak{h})$ and $\Delta_{\mathfrak{l}}=\Delta(\mathfrak{l},\mathfrak{h})$ for the root systems of $\mathfrak{g}$ and $\mathfrak{l}$ with respect to $\mathfrak{h}$, respectively. Write $\Delta(\mathfrak{u})$ for the set of roots of $\mathfrak{h}$ occurring in $\mathfrak{u}$. Fix a positive root system $\Delta_{\mathfrak{g}}^{+}$ containing $\Delta(\mathfrak{u})$. Then $\Delta_{\mathfrak{l}}^{+}:=\Delta_{\mathfrak{l}}\cap\Delta_{\mathfrak{g}}^{+}$ is a positive root system of $\mathfrak{l}$ and $\Delta_{\mathfrak{g}}^{+}=\Delta_{\mathfrak{l}}^{+}\cup\Delta(\mathfrak{u})$. Write $\rho_{\mathfrak{g}}$ and $\rho_{\mathfrak{l}}$ for the half sums of the roots in $\Delta_{\mathfrak{g}}^{+}$ and $\Delta_{\mathfrak{l}}^{+}$, respectively. We will use the notation $\rho(\mathfrak{u})$ for the half sum of the roots occurring in $\mathfrak{u}$. Then $\rho_{\mathfrak{g}}=\rho_{\mathfrak{l}}+\rho(\mathfrak{u})$.
There is an orthogonal decomposition $\mathfrak{g}=\mathfrak{l}\oplus\mathfrak{s}$ with $\mathfrak{s}=\mathfrak{u}\oplus\overline{\mathfrak{u}}$. Note that the restriction of $B$ to $\mathfrak{s}$ is non-degenerate.
Recall the Clifford algebra
${\mathrm{C}}(\mathfrak{s}):={\mathrm{T}}(\mathfrak{s})/(\textrm{the ideal generated by }\{x\otimes x+B(x,x):x\in\mathfrak{s}\}),$
where ${\mathrm{T}}(\mathfrak{s})$ is the tensor algebra over $\mathfrak{s}$. Consequently, ${\mathrm{C}}(\mathfrak{s})$ is generated by $\mathfrak{s}$ with the following defining relations:
(3.1) $x\cdot x=-B(x,x),\quad x\in\mathfrak{s}.$
Since $B$ is $G$-invariant, the adjoint action of $\mathfrak{l}$ on $\mathfrak{s}$ induces a Lie algebra homomorphism
$\mathfrak{l}\rightarrow\mathfrak{s}\mathfrak{o}(\mathfrak{s}).$
Compositing it with the embedding of $\mathfrak{s}\mathfrak{o}(\mathfrak{s})$ into the Clifford algebra ${\mathrm{C}}(\mathfrak{s})$ (cf.  [5, Section 2.1.9]), we obtain a Lie algebra homomorphism
$\nu:\mathfrak{l}\rightarrow{\mathrm{C}}(\mathfrak{s}).$
Then we embed the Lie algebra $\mathfrak{l}$ into ${\mathrm{U}}(\mathfrak{g})\otimes{\mathrm{C}}(\mathfrak{s})$ via
$X\longmapsto X\otimes 1+1\otimes\nu(X),$
where ${\mathrm{U}}$ indicates the universal enveloping algebra. This embedding extends to an algebra homomorphism
(3.2) $\gamma\colon{\mathrm{U}}(\mathfrak{l})\longrightarrow{\mathrm{U}}(\mathfrak{g})\otimes{\mathrm{C}}(\mathfrak{s}).$
We embed the exterior algebra $\wedge^{\bullet}\mathfrak{s}$ into ${\mathrm{T}}(\mathfrak{s})$ as a subspace via
$x_{1}\wedge\cdots\wedge x_{k}\longmapsto\frac{1}{k!}\sum_{\sigma\in S_{k}}\operatorname{sgn}(\sigma)x_{\sigma(1)}\otimes\cdots\otimes x_{\sigma(k)}\quad(k\in\mathbb{N}),$
where $S_{k}$ is the group of permutations on $\{1,\cdots,k\}$, $\operatorname{sgn}(\sigma)$ is the sign of a permutation $\sigma$.
The Chevalley map $\varphi:\wedge^{\bullet}\mathfrak{s}\rightarrow{\mathrm{C}}(\mathfrak{s})$ is obtained by composing this embedding with the natural homomorphism ${\mathrm{T}}(\mathfrak{s})\rightarrow{\mathrm{C}}(\mathfrak{s})$. Let $\{Z_{1},\dots,Z_{n}\}$ be an orthonormal basis of $\mathfrak{s}$. Then the Chevalley map $\varphi$ is the linear map determined by the following formulas:
$\varphi(Z_{i_{1}}\wedge\cdots\wedge Z_{i_{k}})=Z_{i_{1}}\cdots Z_{i_{k}}$
where $1\leq i_{1}<\cdots<i_{k}\leq n.$ The Chevalley map is an isomorphism of vector spaces.
[…]
(3.3) $2\omega(X,Y,Z)=B([X,Y],Z)$
for all $X,Y,Z\in\mathfrak{s}$. To be explicit,
(3.4) $v=\cfrac{1}{2}\sum_{1\leq i<j<k\leq n}B\left([Z_{i},Z_{j}],Z_{k}\right)Z_{i}\wedge Z_{j}\wedge Z_{k}.$
The Kostant’s cubic Dirac operator $D$ with respect to $(\mathfrak{g},\mathfrak{l})$ is the element
$D=\sum_{i=1}^{n}Z_{i}\otimes Z_{i}+1\otimes\varphi(v)\in\mathrm{{\mathrm{U}}}(\mathfrak{g})\otimes\mathrm{{\mathrm{C}}}(\mathfrak{s}).$
The definition of $D$ is independent with the choice of the orthonormal basis.
The following result, which is crucial for our application, is proved by Kostant in [8, Theorem 2.16] in a more general setting.
Theorem 3.1.
The equality
(3.5) $D^{2}=-\Omega_{\mathfrak{g}}\otimes 1+\gamma(\Omega_{\mathfrak{l}})+\textit{C}$
holds in ${\mathrm{U}}(\mathfrak{g})\otimes{\mathrm{C}}(\mathfrak{s})$, where $\Omega_{\mathfrak{g}}$ and $\Omega_{\mathfrak{l}}$ are the Casimir elements for $\mathfrak{g}$ and $\mathfrak{l}$ with respect to $B$, and
$\textit{C}=(\rho_{\mathfrak{l}},\rho_{\mathfrak{l}})-(\rho_{\mathfrak{g}},\rho_{\mathfrak{g}}),$
where $\rho_{\mathfrak{g}},\rho_{\mathfrak{l}}\in\mathfrak{h}^{*}$ are the half sums of positive roots of $\mathfrak{g},\mathfrak{l}$, respectively.
Remark 3.2.
Our definition of Clifford algebra follows the one in [5]. However, Kostant uses a different definition of the Clifford algebra ${\mathrm{C}}(\mathfrak{s})$ in [8], which requires $x\cdot x=B(x,x)$ for $x\in\mathfrak{s}$. Therefore, equalities (3.3) and (3.5) are respectively different from [8, Formula (1.20)] and [8, Formula (2.17)] by some signs.
Since $\mathfrak{s}=\mathfrak{u}\oplus\overline{\mathfrak{u}}$ is even-dimensional, the Clifford algebra ${\mathrm{C}}(\mathfrak{s})$ has a unique irreducible module, which is called the Spin module of ${\mathrm{C}}(\mathfrak{s})$. Fix a nonzero element $\overline{u}_{\mathrm{top}}$ in $\wedge^{\mathrm{top}}\overline{\mathfrak{u}}$, and define a linear isomorphism
(3.6) $\wedge^{\bullet}\mathfrak{u}\longrightarrow\varphi(\wedge^{\bullet}\mathfrak{u}\otimes\wedge^{\mathrm{top}}\overline{\mathfrak{u}})\qquad x\mapsto\varphi(x\otimes\overline{u}_{\mathrm{top}})$
where $\varphi$ is the Chevalley map, and $\wedge^{\bullet}\mathfrak{u}\otimes\wedge^{\mathrm{top}}\overline{\mathfrak{u}}$ is a subspace of $\wedge^{\bullet}\mathfrak{s}=\wedge^{\bullet}\mathfrak{u}\otimes\wedge^{\bullet}\overline{\mathfrak{u}}$. Note that $\varphi(\wedge^{\bullet}\mathfrak{u}\otimes\wedge^{\mathrm{top}}\overline{\mathfrak{u}})$ is the left ideal of ${\mathrm{C}}(\mathfrak{s})$ generated by $\varphi(\wedge^{\mathrm{top}}\overline{\mathfrak{u}})$. Denote by $S$ the vector space $\wedge^{\bullet}\mathfrak{u}$, together with the ${\mathrm{C}}(\mathfrak{s})$-action induced by the linear isomorphism (3.6). Then $S$ is a Spin module of ${\mathrm{C}}(\mathfrak{s})$ (see [5, Section 2.2.2]).
The Killing form (and hence also $B$) induces a non-degenerate pairing $\mathfrak{u}\times\overline{\mathfrak{u}}\to\mathbb{C}$ allowing us to identify the dual of $\overline{\mathfrak{u}}$ with $\mathfrak{u}$ in a canonical way. In particular, for a smooth representation $V$ of $G$, one obtains a canonical degree-preserving, $L$-equivariant isomorphism
(3.7) $\wedge^{\bullet}\mathfrak{u}\otimes V\longrightarrow\wedge^{\bullet}(\overline{\mathfrak{u}}^{*})\otimes V$
between complexes computing ${\mathrm{H}}_{\bullet}(\mathfrak{u},V)$ and ${\mathrm{H}}^{\bullet}(\overline{\mathfrak{u}},V)$, which allows us to consider the coboundary map $\operatorname{d}\,$ on the right hand side and the boundary map $\partial$ on the left hand side as being defined on the same space $\wedge^{\bullet}\mathfrak{u}\otimes V=V\otimes S$.
Theorem 3.3.
[5, Proposition 9.1.6] Under the action of ${\mathrm{U}}(\mathfrak{g})\otimes{\mathrm{C}}(\mathfrak{s})$ on $V\otimes S$, the cubic Dirac operator D acts on $V\otimes S$ as $2\partial+\operatorname{d}\,$.
Theorem 3.4.
Suppose that $V$ is finite-dimensional, the following identities hold:
(3.8) $V\otimes S={\mathrm{Ker}}\,D\oplus{\mathrm{Im}}\,D,$
(3.9) ${\mathrm{Ker}}\,D={\mathrm{Ker}}\,D^{2},\qquad\,\,\,$ ${\mathrm{Im}}\,D={\mathrm{Im}}\,D^{2},$
(3.10) ${\mathrm{Ker}}\,D={\mathrm{Ker}}\operatorname{d}\,\cap\,{\mathrm{Ker}}\,\partial,$ ${\mathrm{Im}}\,D={\mathrm{Im}}\operatorname{d}\,\oplus\,{\mathrm{Im}}\,\partial,$
(3.11) ${\mathrm{Ker}}\operatorname{d}\,={\mathrm{Ker}}\,D\oplus{\mathrm{Im}}\operatorname{d},$ ${\mathrm{Ker}}\,\partial={\mathrm{Ker}}\,D\oplus{\mathrm{Im}}\,\partial.$
All the subspaces of $V\otimes S$ above are $D^{2}$-invariant. Moreover, the restriction of $D^{2}$ to ${\mathrm{Im}}\operatorname{d}$ is bijective and equals $2\!\operatorname{d}\!\partial$. The restriction of $D^{2}$ to ${\mathrm{Im}}\,\partial$ is bijective and equals $2\partial\!\operatorname{d}$.
Proof.
The identities in (3.8) and (3.9) are established in [5, Lemma 9.2.4]. The identities in (3.10) follow from [5, Lemma 9.2.3] and [5, Formula (9.3)]. The identities in (3.11) are proved in [5, Theorem 9.2.5]. Note that the second identity in (3.9) implies that the restriction of $D^{2}$ to ${\mathrm{Im}}\,D$ is bijective. Then the rest of the theorem is easily deduced from the formula $D^{2}=2\partial\!\operatorname{d}\!+2\!\operatorname{d}\!\partial.$
4. Proof of the main theorem
We keep the notations and assumptions from previous sections. In particular, $G$ is a (possibly disconnected) compact Lie group with complexified Lie algebra $\mathfrak{g}$, $(\pi,V)$ is a smooth representation of $G$ and $\mathfrak{q}$ is a parabolic subalgebra of $\mathfrak{g}$ with nilpotent radical $\mathfrak{u}$. Use the notations in Section 3, $S=\wedge^{\bullet}\mathfrak{u}$, $\mathfrak{h}$ is a common Cartan subalgebra of $\mathfrak{l}$ and $\mathfrak{g}$, $D$ is the cubic Dirac operator with respect to $(\mathfrak{g},\mathfrak{l})$ and $\partial\in{\mathrm{End}}(V\otimes S)$ (resp. $\operatorname{d}\in{\mathrm{End}}(V\otimes S)$) are the boundary (resp. coboundary) map of Lie algebra homology (resp. cohomology). From now on, we view $D$ as an operator on $V\otimes S$.
For every $\lambda\in\widehat{G}$, we already fixed a representative $V_{\lambda}$, which has finitely many highest weights as a $\mathfrak{g}$-module with respect to the positive root system $\Delta_{\mathfrak{g}}^{+}$. We choose one of these highest weights and denote it by $\lambda$, by abuse of notation. Similarly, we fix for every $\mu\in\widehat{L}$ in the unitary dual of $L$ a representative $W_{\mu}$ and a highest weight of $W_{\mu}$, which is denoted by $\mu$.
We view $S$ as a representation of $G$ with trivial action, then $V\otimes S$ is also a smooth representation of $G$, and this action of $G$ on $V\otimes S$ is denoted by $\Pi$.
We consider a representation $(\Sigma,V\otimes S)$ of $L$, where the action $\Sigma$ is given by the tensor product of the restriction of $\pi$ and the adjoint action on $S=\wedge^{\bullet}\mathfrak{u}.$ It is worth mentioning that $\operatorname{d}\,,\partial,D$ are $L$-equivariant operators on $(\Sigma,V\otimes S)$.
For $\lambda\in\widehat{G}$ and $\mu\in\widehat{L}$, consider the corresponding continuous linear projectors onto the $\lambda$- and $\mu$-isotypic components of $V\otimes S$:
$P(\lambda):=\Pi(\overline{\chi}_{\lambda}):$ $V\otimes S\longrightarrow V\otimes S,$
$Q(\mu):=\Sigma(\overline{\chi}_{\mu}):$ $V\otimes S\longrightarrow V\otimes S.$
Lemma 4.1.
The element $\Omega_{\mathfrak{g}}\otimes 1\in{\mathrm{U}}(\mathfrak{g})\otimes{\mathrm{C}}(\mathfrak{s})$ acts on ${\mathrm{Im}}\,P(\lambda)$ via the scalar multiplication by
$c(\lambda)=(\lambda+\rho_{\mathfrak{g}},\lambda+\rho_{\mathfrak{g}})-(\rho_{\mathfrak{g}},\rho_{\mathfrak{g}}),$
and $\gamma(\Omega_{\mathfrak{l}})$ acts on ${\mathrm{Im}}\,Q(\mu)$ via the scalar multiplication by
$(\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}},\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}})-(\rho_{\mathfrak{l}},\rho_{\mathfrak{l}}).$
Proof.
It is easy to prove that $\Omega_{\mathfrak{g}}$ acts on $V_{\lambda}$ via the scalar multiplication by
$c(\lambda)=(\lambda+\rho_{\mathfrak{g}},\lambda+\rho_{\mathfrak{g}})-(\rho_{\mathfrak{g}},\rho_{\mathfrak{g}}).$
It follows that the element $\Omega_{\mathfrak{g}}\otimes 1\in{\mathrm{U}}(\mathfrak{g})\otimes{\mathrm{C}}(\mathfrak{s})$ acts on ${\mathrm{Im}}P(\lambda)$ via the scalar multiplication by $c(\lambda)$.
By differentiating the group action $\Sigma$, we obtain the corresponding action of $\mathfrak{l}$ on $V\otimes S$. By abuse of notation, we also denote this action by $\Sigma$. The ${\mathrm{U}}(\mathfrak{g})\otimes{\mathrm{C}}(\mathfrak{s})$-module structure on $V\otimes S$, combined with the morphism $\gamma:{\mathrm{U}}(\mathfrak{l})\rightarrow{\mathrm{U}}(\mathfrak{g})\otimes{\mathrm{C}}(\mathfrak{s})$ (cf. (3.2)), induces another action of $\mathfrak{l}$ on $V\otimes S$. By abuse of notation, we denote this action by $\gamma$. It is proved in [9, Proposisition 3.6] that
$\gamma\cong\Sigma\otimes\rho(\overline{\mathfrak{u}}),$
as $\mathfrak{l}$-modules, where $(\rho(\overline{\mathfrak{u}}),\mathbb{C}_{\rho(\overline{\mathfrak{u}})})$ is the one-dimensional $\mathfrak{l}$-module with weight $\rho(\overline{\mathfrak{u}})=\rho_{\mathfrak{l}}-\rho_{\mathfrak{g}}$.
It suffices to prove that for all $\mu\in\widehat{L}$, the Casimir operator $\Omega_{\mathfrak{l}}$ acts on $W_{\mu}\otimes\mathbb{C}_{\rho(\overline{\mathfrak{u}})}$ via the scalar multiplication by
$(\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}},\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}})-(\rho_{\mathfrak{l}},\rho_{\mathfrak{l}}).$
If $L$ is connected, this result is straightforward. In the general case, a bit more argument is required.
Let $\mathfrak{b}_{L}$ be the Borel subalgebra of $\mathfrak{l}$ associated with $\Delta_{\mathfrak{l}}^{+}$ and $T_{L}:=N_{L}(\mathfrak{b}_{L})$ be the Cartan subgroup of $L$. Recall that $T_{L}$ acts on $\mathfrak{h}^{*}$ by the coadjoint action, which preserves $\Delta_{\mathfrak{l}}^{+}$, and inner form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is invariant under the coadjoint action. By the Cartan-Weyl theory (see [7, Chapter IV, Section 2]), there is a finite set $\{t_{1},\cdots,t_{k}\}$ of elements in the Cartan subgroup $T_{L}$ of $L$, such that as $\mathfrak{l}$-modules,
$W_{\mu}=\bigoplus_{1\leq i\leq k}m_{i}W^{\prime}_{t_{i}.\mu}$
where $W^{\prime}_{t_{i}.\mu}$ is the irreducible $\mathfrak{l}$-module with highest weight $t_{i}.\mu$, and $m_{i}$ denotes the multiplicity of $W^{\prime}_{t_{i}.\mu}$ in $W_{\mu}$. Then $\Omega_{\mathfrak{l}}$ acts on $W^{\prime}_{t_{i}.\mu}\otimes\mathbb{C}_{\rho(\overline{\mathfrak{u}})}$ via the scalar multiplication by
$(t_{i}.\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}},t_{i}.\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}})-(\rho_{\mathfrak{l}},\rho_{\mathfrak{l}}).$
Since $T_{L}$ normalizes $\overline{\mathfrak{u}}$ and $\mathfrak{b}_{L}$, it follows that $\rho(\overline{\mathfrak{u}}),\rho_{\mathfrak{l}}\in\mathfrak{h}^{*}$ are invariant under the coadjoint action of $T_{L}$ on $\mathfrak{h}^{*}$. Therefore,
$(t_{i}.\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}},t_{i}.\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}})-(\rho_{\mathfrak{l}},\rho_{\mathfrak{l}})$
$=$ $(\mu+t_{i}^{-1}.\rho(\overline{\mathfrak{u}})+t_{i}^{-1}.\rho_{\mathfrak{l}},\mu+t_{i}^{-1}.\rho(\overline{\mathfrak{u}})+t_{i}^{-1}.\rho_{\mathfrak{l}})-(\rho_{\mathfrak{l}},\rho_{\mathfrak{l}})$
$=$ $(\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}},\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}})-(\rho_{\mathfrak{l}},\rho_{\mathfrak{l}}).$
This completes the proof of the lemma. ∎
Note that the isotypic component ${\mathrm{Im}}\,P(\lambda)=V(\lambda)\otimes S$ is an invariant subspace for the operators $Q(\mu),\,D,\,\partial$ and $\operatorname{d}$. We denote the restriction of $D,\,\partial$ and $\operatorname{d}$ to endomorphisms of ${\mathrm{Im}}\,P(\lambda)$ by $D_{\lambda},\,\partial_{\lambda},$ and $\operatorname{d}_{\lambda}$, respectively. Then $\partial_{\lambda},\operatorname{d}_{\lambda}\in{\mathrm{End}}(V(\lambda)\otimes S)$ are the boundary and coboundary maps for computing ${\mathrm{H}}_{\bullet}(\mathfrak{u},V(\lambda))$ and ${\mathrm{H}}^{\bullet}(\overline{\mathfrak{u}},V(\lambda))$.
Theorem 4.2.
(a) For all $\lambda\in\widehat{G}$ and $\mu\in\widehat{L}$,
$W(\lambda,\mu):=Q(\mu)\left(V(\lambda)\otimes S\right)$
is a $D^{2}$-invariant closed subspace of $V\otimes S$ on which $D^{2}$ acts via the scalar multiplication by
(4.1) $c(\lambda,\mu):=(\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}},\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}})-(\lambda+\rho_{\mathfrak{g}},\lambda+\rho_{\mathfrak{g}}).$
(b) As representations of $L$:
(4.2) $V(\lambda)\otimes S$ $={\mathrm{Im}}\,D_{\lambda}\oplus{\mathrm{Ker}}\,D_{\lambda},$
(4.3) ${\mathrm{Im}}\,D_{\lambda}$ $={\mathrm{Im}}\,\partial_{\lambda}\oplus{\mathrm{Im}}\operatorname{d}_{\lambda},$
(4.4) ${\mathrm{Ker}}\operatorname{d}_{\lambda}$ $={\mathrm{Im}}\operatorname{d}_{\lambda}\oplus\,{\mathrm{Ker}}\,D_{\lambda},$
(4.5) ${\mathrm{Ker}}\,\partial_{\lambda}$ $={\mathrm{Im}}\,\partial_{\lambda}\oplus{\mathrm{Ker}}\,D_{\lambda},$
(4.6) ${\mathrm{Im}}\,D_{\lambda}$ $={\mathrm{Im}}\,D_{\lambda}^{2}\,=\bigoplus_{\mu\in I(\lambda)}W(\lambda,\mu),$
(4.7) ${\mathrm{Ker}}\,D_{\lambda}$ $={\mathrm{Ker}}\,D_{\lambda}^{2}=\bigoplus_{\mu\in J(\lambda)}W(\lambda,\mu),$
where
$I(\lambda):=\{\mu\in\widehat{L}:{\mathrm{Hom}}_{L}(W_{\mu},V_{\lambda}\otimes S)\not=0\text{ and }c(\lambda,\mu)\not=0\},$
$J(\lambda):=\{\mu\in\widehat{L}:{\mathrm{Hom}}_{L}(W_{\mu},V_{\lambda}\otimes S)\not=0\text{ and }c(\lambda,\mu)=0\}.$
(c) All subspaces of $V\otimes S$ occurring in (4.2)—(4.7) are closed and $D^{2}$-invariant. Moreover, the restriction of $D_{\lambda}^{2}$ to ${\mathrm{Im}}\,\partial_{\lambda}$ is bijective and equals $2\partial_{\lambda}\!\operatorname{d}_{\lambda}$, and the restriction of $D_{\lambda}^{2}$ to ${\mathrm{Im}}\,\operatorname{d}_{\lambda}$ is bijective and equals $2\!\operatorname{d}_{\lambda}\!\partial_{\lambda}$.
Proof.
Combining Theorem 3.1 with Lemma 4.1, we have that $D^{2}$ acts on $W(\lambda,\mu)$ via the scalar multiplication by $c(\lambda,\mu)=(\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}},\mu+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}})-(\lambda+\rho_{\mathfrak{g}},\lambda+\rho_{\mathfrak{g}})$, which in particular shows that $W(\lambda,\mu)$ is stable under $D^{2}$.
We impose ${\mathrm{Hom}}_{G}(V_{\lambda},V)$ with the trivial actions of $G$ and ${\mathrm{U}}(\mathfrak{g})\otimes{\mathrm{C}}(\mathfrak{s})$. It is clear that $V(\lambda)=V_{\lambda}\otimes{\mathrm{Hom}}_{G}(V_{\lambda},V)$ as smooth representations of $G$, which implies that
$V(\lambda)\otimes S=(V_{\lambda}\otimes S)\otimes{\mathrm{Hom}}_{G}(V_{\lambda},V)$
as smooth representations of $G$ and as ${\mathrm{U}}(\mathfrak{g})\otimes{\mathrm{C}}(\mathfrak{s})$-modules. For every subspace $V^{\prime}\subseteq V_{\lambda}\otimes S$ the corresponding subspace $V^{\prime}\otimes{\mathrm{Hom}}_{G}(V_{\lambda},V)$ is closed. Therefore, applying Theorem 3.4 to the factor $V_{\lambda}\otimes S$, applying $-\otimes{\mathrm{Hom}}_{G}(V_{\lambda},V)$ and translating the result back to $V(\lambda)\otimes S$, one obtains the identities (4.2), (4.3), (4.4) and (4.5), as well as the respective first identities in (4.6) and (4.7). Moreover, all spaces which occur are closed in $V\otimes S$.
In order to establish the respective second identities in (4.6) and (4.7), observe that the sets $I(\lambda)$ and $J(\lambda)$ are finite. Consequently,
$V(\lambda)\otimes S=\bigoplus_{\mu\in I(\lambda)\cup J(\lambda)}W(\lambda,\mu)$
is a finite direct sum. The proof of (4.6) and (4.7) is completed by combining this decomposition with the known action of $D_{\lambda}^{2}$ via the scalar $c(\lambda,\mu)$. ∎
Lemma 4.3.
For all $\lambda\in\widehat{G}$,
(4.8) $\sum_{\mu\in I(\lambda)}(\dim W_{\mu})^{2}\leq(\dim V_{\lambda})^{2}(\dim S)^{2}.$
Moreover, there exists a constant $c_{\mathfrak{g}}>0$, only depending on $\mathfrak{g}$, with the property that for all $\lambda\in\widehat{G}$ and all $\mu\in I(\lambda)$:
(4.9) $c_{\mathfrak{g}}\leq\lvert c(\lambda,\mu)\rvert.$
Proof.
In the isotypic decomposition of $V_{\lambda}\otimes S$ as a representation of $L$, the class of $W_{\mu}$ occurs at least once if and only if $\mu\in I(\mu)\cup J(\mu)$. Therefore
$\sum_{\mu\in I(\lambda)}\dim W_{\mu}\leq\dim V_{\lambda}\otimes S.$
Squaring both sides of the inequality and applying the Cauchy-Schwartz inequality yields (4.8).
There is an orthogonal decomposition $\mathfrak{h}^{*}=(\mathfrak{h}^{ss})^{*}\oplus\mathfrak{z}(\mathfrak{g})^{*}$, where $\mathfrak{h}^{ss}:=\mathfrak{h}\cap\mathfrak{g}^{ss}$ and $\mathfrak{z}(\mathfrak{g})$ is the center of $\mathfrak{g}$. Let $\Lambda_{ss}\subseteq(\mathfrak{h}^{ss})^{*}\subseteq\mathfrak{h}^{*}$ denote the lattice of integral weights of $\mathfrak{g}^{ss}$. Define a real-valued function $f$ on $\Lambda_{ss}\times\Lambda_{ss}$ by
$f(x,y)=(y,y)-(x,x),\quad(x,y)\in\Lambda_{ss}\times\Lambda_{ss}.$
Note that $B$ extends the Killing form on $\mathfrak{g}^{ss}$, which implies that (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is rational with respect to the weight lattice $\Lambda_{ss}$. It follows that $f(x,y)$ is a polynomial with rational coefficients on $\Lambda_{ss}\times\Lambda_{ss}$. Consequently, there exists a constant $c_{\mathfrak{g}}>0$ such that $|f(x,y)|\geq c_{\mathfrak{g}}$ whenever $f(x,y)\neq 0$.
For $x\in\mathfrak{h}^{*}$, write $x_{ss}$ and $x_{\mathfrak{z}(\mathfrak{g})}$ for its restrictions to $\mathfrak{h}^{ss}$ and $\mathfrak{z}(\mathfrak{g})$, respectively. It is clear that for all $\lambda\in\widehat{G}$ and all $\mu\in I(\lambda)$,
$\lambda_{\mathfrak{z}(\mathfrak{g})}=\mu_{\mathfrak{z}(\mathfrak{g})}\quad\text{and}\quad c(\lambda,\mu)=c(\lambda_{ss},\mu_{ss})=f(\lambda_{ss}+\rho_{\mathfrak{g}},\mu_{ss}+\rho(\overline{\mathfrak{u}})+\rho_{\mathfrak{l}})\neq 0.$
Consequently, $\lvert c(\lambda,\mu)\rvert\geq c_{\mathfrak{g}}$ for all $\lambda\in\widehat{G}$ and all $\mu\in I(\lambda)$. ∎
For all $\lambda\in\widehat{G}$, by decomposition (4.6), one knows that $D_{\lambda}^{2}$ induces an (algebraic) automorphism of ${\mathrm{Im}}\,D_{\lambda}$. Using decomposition (4.2), let $T_{\lambda}$ be the linear operator on $V(\lambda)\otimes S$, which is inverse of $D_{\lambda}^{2}$ on ${\mathrm{Im}}\,D_{\lambda}$ and acts as the zero operator on ${\mathrm{Ker}}D_{\lambda}$. To be explicit, the formula of $T_{\lambda}$ is given by
(4.10) $T_{\lambda}v(\lambda)=\sum_{\mu\in I(\lambda)}\frac{1}{c(\lambda,\mu)}Q(\mu)v(\lambda)$
for all $v(\lambda)\in V(\lambda)\otimes S$. It is clear that $T_{\lambda}$ is $L$-equivariant and
(4.11) $D^{2}_{\lambda}T_{\lambda}=T_{\lambda}D^{2}_{\lambda}=(T_{\lambda}D^{2}_{\lambda})^{2}.$
Lemma 4.4.
For every continuous seminorm $\lvert\,\cdot\,\rvert_{p}$ on $V\otimes S$, there is a continuous seminorm $\lvert\,\cdot\,\rvert_{q}$ such that
(4.12) $\lvert T_{\lambda}v(\lambda)\rvert_{p}\leq c_{\mathfrak{g}}^{-1}(\dim S)^{2}(\dim V_{\lambda})^{2}\lvert v(\lambda)\rvert_{q}$
for all $\lambda\in\widehat{G}$, $v(\lambda)\in V(\lambda)\otimes S$, where $c_{\mathfrak{g}}>0$ is the constant in Lemma 4.3. In particular, $T_{\lambda}$ is continuous.
Proof.
Applying the bound (4.9) from Lemma 4.3 to (4.10), we obtain
(4.13) $\lvert T_{\lambda}v(\lambda)\rvert_{p}\leq c_{\mathfrak{g}}^{-1}\sum_{\mu\in I(\lambda)}\lvert Q(\mu)v(\lambda)\rvert_{p}$
for all $v(\lambda)\in V(\lambda)\otimes S$.
Invoking Lemma 2.3 for the representation $(\Sigma,V\otimes S)$, there is a continuous seminorm $\lvert\,\cdot\,\rvert_{q}$ on $V\otimes S$ such that
$\lvert Q(\mu)v\rvert_{p}\leq(\dim W_{\mu})^{2}\lvert v\rvert_{q}$
for all $v\in V\otimes S$.
Summing over all $\mu\in I(\lambda)$ and applying the bound (4.8) from Lemma 4.3, we obtain
(4.14) $\sum_{\mu\in I(\lambda)}\lvert Q(\mu)v(\lambda)\rvert_{p}\leq(\dim V_{\lambda})^{2}(\dim S)^{2}\lvert v(\lambda)\rvert_{q}$
for all $v(\lambda)\in V(\lambda)\otimes S$.
Combining (4.13) and (4.14), we obtain
(4.15) $\lvert T_{\lambda}v(\lambda)\rvert_{p}\leq c_{\mathfrak{g}}^{-1}(\dim V_{\lambda})^{2}(\dim S)^{2}\lvert v(\lambda)\rvert_{q}.$
Therefore, $T_{\lambda}$ is continuous. ∎
We augment the family of operators $T_{\lambda}$ for $\lambda\in\widehat{G}$ to an operator on $V\otimes S$.
Lemma 4.5.
There is a unique $L$-equivariant continuous linear operator $T$ on $V\otimes S$ whose restriction to $V(\lambda)\otimes S$ is $T_{\lambda}$ for every $\lambda\in\widehat{G}$. Moreover, $T$ commutes with $D^{2}$, and $D^{2}T$ is an $L$-equivariant continuous linear projection.
Proof.
We aim at defining the $L$-equivariant operator $T$ by the formula
(4.16) $Tv=\sum_{\lambda\in\widehat{G}}T_{\lambda}P(\lambda)v,\quad\text{for~{}}v\in V\otimes S.$
For if $T$ exists, it must be of this form. We need to show that $T$ is well defined. That is the right hand side is absolutely convergent for every $v\in V\otimes S$ and that the resulting operator $T$ is continuous.
Let $\lvert\,\cdot\,\rvert_{p}$ be a continuous seminorm on $V\otimes S$, by Lemma 4.4, there is a continuous seminorm $\lvert\,\cdot\,\rvert_{q}$ such that for all $v\in V\otimes S$,
$\lvert T_{\lambda}P(\lambda)v\rvert_{p}\leq c_{\mathfrak{g}}^{-1}(\dim S)^{2}(\dim V_{\lambda})^{2}\lvert P(\lambda)v\rvert_{q}$
where $c_{\mathfrak{g}}$ is a positive constant depending only on $\mathfrak{g}$.
Using Lemma 2.3, there is a continuous seminorm $\lvert\,\cdot\,\rvert_{r}$ such that for all $m\in{\mathbb{N}}$ and all $v\in V\otimes S$,
$\lvert P(\lambda)v\rvert_{q}\leq(1+c(\lambda))^{-m}(\dim V_{\lambda})^{2}\lvert(1+\Omega_{\mathfrak{g}})^{m}v\rvert_{r}.$
Hence, for all $m\in{\mathbb{N}}$ and all $v\in V\otimes S$
$\sum_{\lambda\in\widehat{G}}\lvert T_{\lambda}P(\lambda)v\rvert_{p}$$\leq$ $c_{\mathfrak{g}}^{-1}\sum_{\lambda\in\widehat{G}}(\dim S)^{2}(\dim V_{\lambda})^{2}\lvert P(\lambda)v\rvert_{q}$
$\leq$ $\left(\sum_{\lambda\in\widehat{G}}(\dim V_{\lambda})^{4}(1+c(\lambda))^{-m}\right)c_{\mathfrak{g}}^{-1}\cdot(\dim S)^{2}\cdot\lvert(1+\Omega_{\mathfrak{g}})^{m}v\rvert_{r}.$
By Lemma 2.4, there is an $m_{0}\in{\mathbb{N}}$ such that
$C_{0}:=c_{\mathfrak{g}}^{-1}\cdot(\dim S)^{2}\sum_{\lambda\in\widehat{G}}(\dim V_{\lambda})^{4}(1+c(\lambda))^{-m}<\infty.$
We conclude that
(4.17) $\sum_{\lambda\in\widehat{G}}\lvert T_{\lambda}P(\lambda)v\rvert_{p}\leq C_{0}\lvert(1+\Omega_{\mathfrak{g}})^{m_{0}}v\rvert_{r}.$
Since $\lvert\,\cdot\,\rvert_{p}$ is an arbitrary continuous seminorm on $V\otimes S$, the Fourier series
$\sum_{\lambda\in\widehat{G}}T_{\lambda}P(\lambda)v$
is absolutely convergent and (4.16) defines a linear operator $T$ on $V\otimes S$. Moreover, for every continuous seminorm $\lvert\,\cdot\,\rvert_{p}$, we have
(4.18) $\lvert Tv\rvert_{p}\leq\sum_{\lambda\in\widehat{G}}\lvert T_{\lambda}P(\lambda)v\rvert_{p}\leq C_{0}\lvert(1+\Omega_{\mathfrak{g}})^{m_{0}}v\rvert_{r}=C_{0}\lvert v\rvert_{t},$
where $\lvert\,\cdot\,\rvert_{t}$ is the continuous seminorm given by $\lvert v\rvert_{t}=\lvert(1+\Omega_{\mathfrak{g}})^{m_{0}}v\rvert_{r}$. This establishes the continuity of $T$.
By construction of $T$ (cf. (4.16)), the relations $D_{\lambda}^{2}T_{\lambda}=T_{\lambda}D_{\lambda}^{2}=(T_{\lambda}D_{\lambda}^{2})^{2}$ (cf. (4.11)) imply the relations $D^{2}T=TD^{2}=(TD^{2})^{2}$. ∎
With the $L$-equivariant continuous projection $D^{2}T$ at hand, we now proceed to prove our main theorem.
Theorem 4.6.
Let $G$ be a compact Lie group and $V$ be a smooth representation of $G$. Then subspaces ${\mathrm{Im}}\,D,{\mathrm{Im}}\,\partial,{\mathrm{Im}}\operatorname{d}\!$ of $V\otimes S$ are closed. Moreover, the following decompositions of smooth representations of $L$ hold:
(4.19) $V\otimes S$ $=$ ${\mathrm{Ker}}\,D\oplus{\mathrm{Im}}\,D,$
(4.20) ${\mathrm{Ker}}\,\partial$ $=$ ${\mathrm{Ker}}\,D\oplus{\mathrm{Im}}\,\partial,$
(4.21) ${\mathrm{Ker}}\operatorname{d}$ $=$ ${\mathrm{Ker}}\,D\oplus{\mathrm{Im}}\operatorname{d}\!.$
Therefore, ${\mathrm{H}}_{\bullet}(\mathfrak{u},V)$ and ${\mathrm{H}}^{\bullet}(\overline{\mathfrak{u}},V)$ are Hausdorff and quasi-complete. Moreover,
(4.22) ${\mathrm{H}}_{\bullet}(\mathfrak{u},V)\cong{\mathrm{H}}^{\bullet}(\overline{\mathfrak{u}},V)\cong{\mathrm{Ker}}\,D$
as smooth representations of $L$.
Proof.
By Lemma 4.5, $D^{2}T$ is an $L$-equivariant continuous projection, therefore the image and kernel of $D^{2}T$ are closed and we have a decomposition of smooth representations of $L$:
$V\otimes S={\mathrm{Ker}}\,D^{2}T\oplus{\mathrm{Im}}\,D^{2}T.$
On the one hand, Lemma 2.2 together with ${\mathrm{Im}}\,D_{\lambda}^{2}T_{\lambda}={\mathrm{Im}}\,D_{\lambda}^{2}={\mathrm{Im}}\,D_{\lambda}$ (cf. (4.6) and (4.11)) shows that
${\mathrm{Im}}\,D^{2}T=\overline{\bigoplus_{\lambda\in\widehat{G}}{\mathrm{Im}}\,D_{\lambda}^{2}T_{\lambda}}=\overline{\bigoplus_{\lambda\in\widehat{G}}{\mathrm{Im}}\,D_{\lambda}}=\overline{{\mathrm{Im}}\,D}.$
[…]
${\mathrm{Im}}\,D^{2}T\subset{\mathrm{Im}}\,D^{2}\subset{\mathrm{Im}}\,D\subset\overline{{\mathrm{Im}}\,D}={\mathrm{Im}}\,D^{2}T.$
We conclude that ${\mathrm{Im}}\,D={\mathrm{Im}}\,D^{2}T$ is closed in $V\otimes S$.
Combining Theorem 4.2 with (4.11), it follows that
${\mathrm{Ker}}\,D_{\lambda}={\mathrm{Ker}}\,D_{\lambda}^{2}={\mathrm{Ker}}\,D^{2}_{\lambda}T_{\lambda}={\mathrm{Ker}}\operatorname{d}_{\lambda}\cap\,{\mathrm{Ker}}\,\partial_{\lambda}.$
Applying Lemma 2.2 once more, we conclude that
(4.23) ${\mathrm{Ker}}\,D={\mathrm{Ker}}\,D^{2}T={\mathrm{Ker}}\operatorname{d}\cap\,{\mathrm{Ker}}\,\partial.$
Therefore, we obtain the decomposition (4.19) of smooth representations of $L$ as claimed.
Note that $\partial_{\lambda},\operatorname{d}_{\lambda}$ commute with $T_{\lambda},D^{2}_{\lambda}$ for all $\lambda\in\widehat{G}$. It follows that $\partial,\operatorname{d}$ commute with $T,D^{2}$. Then ${\mathrm{Ker}}\,\partial$ and ${\mathrm{Ker}}\operatorname{d}$ are closed subspaces which are invariant under $D^{2}$ and $T$. In addition, ${\mathrm{Ker}}\,\partial$ and ${\mathrm{Ker}}\operatorname{d}$ are $L$-invariant subspaces. Therefore $D^{2}T$ restricted to ${\mathrm{Ker}}\,\partial$ and ${\mathrm{Ker}}\operatorname{d}\,$ is again an $L$-equivariant continuous projection operator on ${\mathrm{Ker}}\,\partial$ and ${\mathrm{Ker}}\operatorname{d}$. Combining this fact with (4.23) shows that
${\mathrm{Ker}}\,\partial$ $=$ ${\mathrm{Ker}}\,(D^{2}T|_{{\mathrm{Ker}}\,\partial})\oplus{\mathrm{Im}}\,(D^{2}T|_{{\mathrm{Ker}}\,\partial})={\mathrm{Ker}}\,D\oplus D^{2}T({\mathrm{Ker}}\,\partial),$
${\mathrm{Ker}}\operatorname{d}$ $=$ ${\mathrm{Ker}}\,(D^{2}T|_{{\mathrm{Ker}}\operatorname{d}})\oplus{\mathrm{Im}}\,(D^{2}T|_{{\mathrm{Ker}}\operatorname{d}})={\mathrm{Ker}}\,D\oplus D^{2}T({\mathrm{Ker}}\operatorname{d}),$
as smooth representations of $L$.
By Theorem 4.2, identities (4.4) (4.5) therein and Lemma 2.2, we have
${\mathrm{Im}}\,(D^{2}T|_{{\mathrm{Ker}}\,\partial})$ $=$ $D^{2}T({\mathrm{Ker}}\,\partial)=\overline{\bigoplus_{\lambda\in\widehat{G}}D^{2}_{\lambda}T_{\lambda}({\mathrm{Ker}}\,\partial_{\lambda})}=\overline{\bigoplus_{\lambda\in\widehat{G}}{\mathrm{Im}}\,\partial_{\lambda}\,}=\overline{{\mathrm{Im}}\,\partial},$
${\mathrm{Im}}\,(D^{2}T|_{{\mathrm{Ker}}\operatorname{d}})$ $=$ $D^{2}T({\mathrm{Ker}}\operatorname{d})=\overline{\bigoplus_{\lambda\in\widehat{G}}D^{2}_{\lambda}T_{\lambda}({\mathrm{Ker}}\operatorname{d}_{\lambda})}=\overline{\bigoplus_{\lambda\in\widehat{G}}{\mathrm{Im}}\operatorname{d}_{\lambda}}=\overline{{\mathrm{Im}}\operatorname{d}}.$
By Theorem 3.3, we have $D^{2}=2\partial\!\operatorname{d}+2\!\operatorname{d}\!\partial$, which in turn implies
$D^{2}|_{{\mathrm{Ker}}\,\partial}=2\partial\!\operatorname{d}\!|_{{\mathrm{Ker}}\,\partial}\quad\text{and}\quad D^{2}|_{{\mathrm{Ker}}\operatorname{d}}=2\!\operatorname{d}\!\partial|_{{\mathrm{Ker}}\operatorname{d}}.$
Consequently, there are chains of inclusions
${\mathrm{Im}}\,\partial\subset\overline{{\mathrm{Im}}\,\partial}\,=D^{2}T({\mathrm{Ker}}\,\partial)=2\partial\!\operatorname{d}T({\mathrm{Ker}}\,\partial)\subset{\mathrm{Im}}\,\partial,$
${\mathrm{Im}}\operatorname{d}\,\subset\overline{{\mathrm{Im}}\operatorname{d}}=D^{2}T({\mathrm{Ker}}\operatorname{d})\,=2\!\operatorname{d}\!\partial T({\mathrm{Ker}}\operatorname{d})\subset{\mathrm{Im}}\operatorname{d}\!.$
This proves decompositions (4.20) and (4.21) of smooth representations of $L$. ∎